### 02 Apr 2026 v1

# A Digit-Sum Formula for Böttcher Coordinates

Abstract.
Let $p$ be an odd prime and let $\phi(x)=x^{p^{2}}+p^{2}x^{p^{2}+1}$. We prove an unconditional digit-sum congruence formula for the coefficients of the Böttcher coordinate of $\phi$. The argument separates the divisible case $k\equiv 0\pmod{p}$ from the general case, gives a complete vector-partition description of the surviving $B$-term monomials, and reduces the remaining partition sum to a one-variable cumulant. As consequences we recover the residue-class congruences conjectured by Salerno and Silverman and obtain an explicit description of the first block of coefficients modulo $p$.
Key words and phrases:
Böttcher coordinate, superattracting germ, non-archimedean dynamics
2020 Mathematics Subject Classification:
37P35, 37F50, 11S82
1. Introduction
The local study of holomorphic iteration begins with the functional equations of Schröder and Koenigs, and the theory of complex dynamics took its modern shape through the early twentieth-century work of Böttcher, Fatou, and Julia [1, 6]. For a one-dimensional analytic germ with a superattracting fixed point, Böttcher’s theorem provides a canonical local coordinate in which the map becomes a pure monomial. In polynomial dynamics this coordinate at infinity does considerably more than linearize the map: it identifies the basin of infinity with the exterior disk, encodes the escape-rate function, and underlies external rays, landing theory, and the combinatorial study of Julia sets and the Mandelbrot set [2, 3, 6]. The same normal form also appears in several variables, where the local geometry near superattracting fixed points becomes subtler [5].
In recent years Böttcher coordinates have also become important in non-archimedean dynamics. Salerno and Silverman studied their integrality properties for one-dimensional superattracting germs over $p$-adic rings [8], while Fu and Nie later settled a conjecture of Salerno–Silverman on the radius of convergence in the wild superattracting case [4]. The family relevant for the present paper is
$\phi(x)=x^{p^{2}}+p^{2}x^{p^{2}+1}.$
If
$f_{\phi}(x)=x\sum_{k\geq 0}\frac{a_{k}}{k!}x^{k}$
denotes the Böttcher coordinate characterized by
$\phi\bigl(f_{\phi}(x)\bigr)=f_{\phi}(x^{p^{2}}),$
then Salerno and Silverman recorded striking mod-$p$ patterns in the coefficients $a_{k}$; their original formulation is Conjecture 25 of [8]. For an integer $n=\sum_{i\geq 0}n_{i}p^{i}$ with $0\leq n_{i}\leq p-1$, let
$S_{p}(n):=\sum_{i\geq 0}n_{i}$
be the sum of its base-$p$ digits.
From now on, we assume that $p\geq 3$. We note that the corresponding result for $p=2$ can be established similarly, but since the arguments for all $p\geq 3$ fit into a unified framework, we do not treat the case $p=2$ in the present paper. The following are our main theorems.
Theorem 1.1 (Digit-sum formula).
Let $k\geq 1$ and write
$k=a+d_{1}p+d_{2}p^{2}+\cdots+d_{r}p^{r},\qquad 0\leq a,d_{i}\leq p-1,$
with
$s:=d_{1}+\cdots+d_{r}=S_{p}(k)-a.$
Then
(1.1) $a_{k}\equiv(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
In fact, we have a stronger result
(1.2) $a_{k}\equiv(-1)^{S_{p}(k)}(a+1)^{S_{p}(k)-1}\pmod{p}$.
We also recover the original residue-class conjecture of Salerno and Silverman.
Theorem 1.2 (Salerno–Silverman’s Conjecture 25).
For every integer $m\geq 1$ one has
$a_{pm}\equiv(-1)^{m},\qquad a_{pm-1}\equiv 0,\qquad a_{pm-2}\equiv-1\pmod{p}$.
The proof is organized by separating the residue class $a=0$ from the classes $a\neq 0$. Section 2 recalls the AABBCC recursion and proves the basic $A$- and $C$-term lemmas. Section 3 supplies a complete proof of the divisible case $a=0$. The remaining sections treat $a\neq 0$: Section 4 gives a vector-partition expansion of the $B$-term, Section 5 evaluates the first block of coefficients and collapses the multivariate cumulant, and Section 6 proves (1.1). In the final section we prove (1.2) and deduce Theorem 1.2.
2. The recursion and two structural lemmas
Let $q:=p^{2}$. As in [8, §5], write
$f_{\phi}(x)=x\sum_{k\geq 0}\frac{a_{k}}{k!}x^{k},\qquad a_{0}=1,$
so that
$\phi\bigl(f_{\phi}(x)\bigr)=f_{\phi}(x^{q}).$
The coefficient comparison may be organized in the form
(2.1) $a_{k}=A_{k}[x^{k}]-B_{k}[x^{k}]-C_{k}[x^{k}],$
where
$A_{k}:=\frac{k!}{q}\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{q\ell},\qquad B_{k}:=\frac{k!}{q}\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{q},\qquad C_{k}:=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{q+1}.$
Here $P[x^{n}]$ denotes the coefficient of $x^{n}$ in the power series $P(x)$. Throughout we use Legendre’s formula
(2.2) $\operatorname{ord}_{p}(n!)=\frac{n-S_{p}(n)}{p-1}$
and the standard digit-sum subadditivity
(2.3) $S_{p}(u+v)\leq S_{p}(u)+S_{p}(v)$
for nonnegative integers $u,v$, together with its obvious extension to finite sums.
Proposition 2.1 (The $A$-term always vanishes modulo $p$).
For every $k\geq 1$ one has
$A_{k}[x^{k}]\equiv 0\pmod{p}$.
Proof.
If $q\nmid k$, then $A_{k}[x^{k}]=0$ because every monomial occurring in $A_{k}$ has degree divisible by $q=p^{2}$.
Assume now that $k=qn=p^{2}n$. Then only the term $\ell=n$ contributes, so
$A_{k}[x^{k}]=\frac{(qn)!}{q\,n!}\,a_{n}=\frac{(qn)!}{p^{2}n!}\,a_{n}.$
Since multiplying by $p^{2}$ only shifts the base-$p$ digits, we have $S_{p}(qn)=S_{p}(n)$. Hence
$\operatorname{ord}_{p}\!\left(\frac{(qn)!}{p^{2}n!}\right)$ $=\operatorname{ord}_{p}((qn)!)-2-\operatorname{ord}_{p}(n!)$
$=\frac{qn-S_{p}(qn)}{p-1}-2-\frac{n-S_{p}(n)}{p-1}$
$=\frac{(p^{2}-1)n}{p-1}-2$
$=(p+1)n-2.$
Since $n\geq 1$ and $p\geq 3$, this is at least $1$. Therefore $p\mid A_{k}[x^{k}]$, so $A_{k}[x^{k}]\equiv 0\pmod{p}$. ∎
For the $B$-term, let $a^{e}=\prod_{n\geq 0}a_{n}^{e_{n}}$ be a monomial occurring in $B_{k}[x^{k}]$. Then
$\sum_{n\geq 0}e_{n}=q,\qquad\sum_{n\geq 0}ne_{n}=k,$
and the coefficient of $a^{e}$ in $B_{k}[x^{k}]$ is
$B_{k}[x^{k}][a^{e}]=\frac{k!}{q}\binom{q}{e_{0},e_{1},\dots}\prod_{n\geq 0}\frac{1}{(n!)^{e_{n}}}.$
A direct application of Legendre’s formula yields
(2.4) $(p-1)\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=-2p+1-S_{p}(k)+\sum_{n\geq 0}S_{p}(e_{n})+\sum_{n\geq 0}e_{n}S_{p}(n).$
Equivalently, with
$c(e):=\frac{\sum_{n\geq 0}e_{n}S_{p}(n)-S_{p}(k)}{p-1},\qquad d(e):=\frac{\sum_{n\geq 0}S_{p}(e_{n})-1}{p-1},$
one has
(2.5) $\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=c(e)+d(e)-2.$
The two quantities $c(e)$ and $d(e)$ measure the carry defects in the base-$p$ additions $\sum_{n}e_{n}n=k$ and $\sum_{n}e_{n}=p^{2}$, respectively.
Proposition 2.2 (A global $C$-term lemma).
If $k\equiv a\pmod{p}$ with $1\leq a\leq p-1$, then
$C_{k}[x^{k}]\equiv a\,a_{k-1}\pmod{p}$.
If $p\mid k$, then
$C_{k}[x^{k}]\equiv 0\pmod{p}$.
Proof.
Let $a^{e}=\prod_{n=0}^{k-1}a_{n}^{e_{n}}$ be a monomial occurring in $C_{k}[x^{k}]$. Since
$C_{k}=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{p^{2}+1},$
its monomials satisfy
$\sum_{n=0}^{k-1}e_{n}=p^{2}+1,\qquad\sum_{n=0}^{k-1}ne_{n}=k-1.$
The coefficient of $a^{e}$ in $C_{k}[x^{k}]$ is
$C_{k}[x^{k}][a^{e}]=k!\binom{p^{2}+1}{e_{0},e_{1},\dots,e_{k-1}}\prod_{n=0}^{k-1}\frac{1}{(n!)^{e_{n}}},$
and a direct Legendre computation gives
(2.6) $(p-1)\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)=-S_{p}(k)-1+\Sigma(e)+T(e),$
where
$\Sigma(e):=\sum_{n=0}^{k-1}S_{p}(e_{n}),\qquad T(e):=\sum_{n=0}^{k-1}e_{n}S_{p}(n).$
Assume first that $p\mid k$. Write $k=p^{\nu}u$ with $\nu\geq 1$ and $p\nmid u$. Subtracting $1$ from $k$ changes the final $\nu$ zeros in the base-$p$ expansion into p1p-1’s and lowers the next digit by $1$, so
$S_{p}(k-1)=S_{p}(k)-1+\nu(p-1).$
Also,
$\Sigma(e)\geq S_{p}(p^{2}+1)=2,\qquad T(e)\geq S_{p}(k-1).$
Therefore every monomial coefficient in $C_{k}[x^{k}]$ satisfies
$(p-1)\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)$ $\geq-S_{p}(k)-1+2+S_{p}(k-1)$
$=\nu(p-1)\geq p-1.$
Hence $\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)\geq 1$ for every monomial, and so $C_{k}[x^{k}]\equiv 0\pmod{p}$.
Now assume that $k\equiv a\pmod{p}$ with $1\leq a\leq p-1$. Then
$S_{p}(k-1)=S_{p}(k)-1.$
Since again $\Sigma(e)\geq 2$ and $T(e)\geq S_{p}(k-1)$, every monomial that survives modulo $p$ must satisfy
(2.7) $\Sigma(e)=2,\qquad T(e)=S_{p}(k-1).$
The condition $\Sigma(e)=2$ leaves exactly two multiplicity patterns:
$e_{t}=p^{2}+1\qquad\text{or}\qquad e_{r}=p^{2},\ e_{s}=1,$
with all other multiplicities equal to 0.
In the first case,
$k-1=(p^{2}+1)t,\qquad T(e)=(p^{2}+1)S_{p}(t).$
If $t>0$, then
$S_{p}(k-1)=S_{p}((p^{2}+1)t)\leq S_{p}(p^{2}t)+S_{p}(t)=2S_{p}(t)<(p^{2}+1)S_{p}(t)=T(e),$
contradicting (2.7). Hence $t=0$, so $k=1$, and the claimed congruence is immediate.
In the second case,
$k-1=p^{2}r+s,\qquad T(e)=p^{2}S_{p}(r)+S_{p}(s).$
If $r>0$, then
$S_{p}(k-1)=S_{p}(p^{2}r+s)\leq S_{p}(r)+S_{p}(s)<p^{2}S_{p}(r)+S_{p}(s)=T(e),$
again a contradiction. Therefore $r=0$, so s=k1s=k-1 and the unique surviving monomial is
$e_{0}=p^{2},\qquad e_{k-1}=1.$
Its coefficient is
$k!\binom{p^{2}+1}{p^{2},1}\frac{1}{(k-1)!}=k(p^{2}+1)\equiv k\equiv a\pmod{p}$.
Thus
$C_{k}[x^{k}]\equiv a\,a_{k-1}\pmod{p}$,
as claimed. ∎
Theorem 2.3 (The case $a=0$).
For every integer $M\geq 1$ one has
$a_{pM}\equiv(-1)^{M}\pmod{p}$.
The proof of Theorem 2.3 is given in the next section.
3. The Case $a=0$
This section gives a complete proof of Theorem 2.3. The main new input is an explicit coefficient identity for a truncated exponential series.
Lemma 3.1 (A truncated-exponential coefficient identity).
For $k\geq 1$, define
$S_{k-1}(y):=\sum_{\ell=0}^{k-1}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
Then
(3.1) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=[y^{k}]e^{-qy/p}-q\frac{(-1)^{k}}{p^{k}k!}.$
Equivalently,
(3.2) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=\frac{(-q/p)^{k}}{k!}-q\frac{(-1)^{k}}{p^{k}k!}.$
Since $q=p^{2}$, this becomes
(3.3) $y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=\frac{(-p)^{k}}{k!}-q\frac{(-1)^{k}}{p^{k}k!}.$
Proof.
From the exponential series,
$e^{-y/p}=\sum_{\ell=0}^{\infty}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
Hence
$S_{k-1}(y)=e^{-y/p}-R_{k}(y),\qquad R_{k}(y):=\sum_{\ell=k}^{\infty}\frac{(-1)^{\ell}}{p^{\ell}\ell!}y^{\ell}.$
In particular,
$R_{k}(y)=\frac{(-1)^{k}}{p^{k}k!}y^{k}+O(y^{k+1}),$
so
$S_{k-1}(y)=E(y)+U(y),\qquad E(y):=e^{-y/p},\qquad U(y):=-\frac{(-1)^{k}}{p^{k}k!}y^{k}+O(y^{k+1}).$
Now expand the $q$th power:
$(E+U)^{q}=E^{q}+qE^{q-1}U+\sum_{j=2}^{q}\binom{q}{j}E^{q-j}U^{j}.$
Since $\operatorname{ord}_{y}(U)\geq k$, every term with $j\geq 2$ has $y$-adic order at least $2k>k$, hence contributes nothing to the coefficient of $y^{k}$. Therefore
$y^{k}](E+U)^{q}=[y^{k}]E^{q}+[y^{k}](qE^{q-1}U).$
Because $E(0)=1$, we have $E^{q-1}(y)=1+O(y)$, and hence
$y^{k}](E^{q-1}U)=[y^{k}]U=-\frac{(-1)^{k}}{p^{k}k!}.$
Substituting gives
$y^{k}]\bigl(S_{k-1}(y)\bigr)^{q}=[y^{k}]E(y)^{q}-q\frac{(-1)^{k}}{p^{k}k!}.$
Finally, $E(y)^{q}=e^{-qy/p}$, which proves (3.1); equation (3.2) follows immediately from the exponential series. ∎
Corollary 3.2.
For every $m\geq 1$,
$y^{m}]\left(\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}\right)^{q}=\frac{p^{m}}{m!}-\frac{q}{p^{m}m!}.$
Proof.
Apply Lemma 3.1 with $k=m$ and replace $y$ by y-y. Since
$\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}=S_{m-1}(-y),$
we obtain
$y^{m}]\left(\sum_{i=0}^{m-1}\frac{y^{i}}{p^{i}i!}\right)^{q}=[y^{m}]e^{qy/p}-q\frac{1}{p^{m}m!}.$
Because $q=p^{2}$, we have $e^{qy/p}=e^{py}$, whose $y^{m}$-coefficient is $p^{m}/m!$. ∎
Proposition 3.3 (Structure of the surviving $B$-term monomials).
Assume that $p\mid k$. Write
$B_{k}[x^{k}]=\sum_{e}\gamma_{B}(e)a^{e},$
where the sum is over all multi-indices $e=(e_{0},\dots,e_{k-1})$ satisfying
$\sigma(e):=\sum_{n=0}^{k-1}e_{n}=q,\qquad\nu(e):=\sum_{n=1}^{k-1}ne_{n}=k,\qquad a^{e}:=\prod_{n=0}^{k-1}a_{n}^{e_{n}}.$
Suppose that a coefficient $\gamma_{B}(e)$ has $p$-adic valuation 0. Then either
$a^{e}=a_{0}^{q-p}a_{1}^{p},$
or else every positive index $n$ with $e_{n}>0$ is divisible by $p$.
Proof.
For such $e$, the coefficient of $a^{e}$ in $B_{k}[x^{k}]$ is
$\gamma_{B}(e):=\frac{k!}{q}\binom{q}{e_{0},\dots,e_{k-1}}\prod_{n=0}^{k-1}\frac{1}{(n!)^{e_{n}}}.$
A straightforward calculation with Legendre’s formula gives
(3.4) $(p-1)\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=-2p+1-S_{p}(k)+\sum_{n=0}^{k-1}\bigl(S_{p}(e_{n})+e_{n}S_{p}(n)\bigr).$
It is convenient to write
$c(e):=\frac{\sum_{n=1}^{k-1}e_{n}S_{p}(n)-S_{p}(k)}{p-1},\qquad d(e):=\frac{\sum_{n=0}^{k-1}S_{p}(e_{n})-1}{p-1}.$
By digit-sum subadditivity, both numerators are nonnegative; since $S_{p}(N)\equiv N\pmod{p-1}$ for every integer $N$, both numerators are divisible by p1p-1, so $c(e),d(e)\in\mathbf{Z}_{\geq 0}$. Using
$\sum_{n=0}^{k-1}e_{n}=q=p^{2},\qquad S_{p}(q)=1,$
we can rewrite (3.4) as
(3.5) $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)+d(e)-2.$
Now assume that $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=0$.
Step 1: if some occurring positive index is not divisible by $p$, then $c(e)\geq 1$. Suppose there exists $n\geq 1$ with $e_{n}>0$ and $p\nmid n$. In the base-$p$ sum
$\sum_{n=1}^{k-1}e_{n}n=k,$
the units digit on the left receives a nonzero contribution from such an $n$, while the right-hand side has units digit 0 because $p\mid k$. Therefore at least one carry must occur from the units place to the $p$-place. Equivalently, the digit-sum defect
$\sum_{n=1}^{k-1}e_{n}S_{p}(n)-S_{p}(k)$
is positive, so $c(e)\geq 1$.
Step 2: $d(e)\neq 0$. Assume for contradiction that $d(e)=0$. Then
$\sum_{n=0}^{k-1}S_{p}(e_{n})=1.$
Since every $e_{n}\geq 0$ and $\sum e_{n}=q=p^{2}$, the only possibility is that one index $m$ satisfies $e_{m}=p^{2}$ and all other $e_{n}$ vanish. In that case,
$k=\sum_{n=1}^{k-1}ne_{n}=mp^{2},$
and therefore
$c(e)=\frac{p^{2}S_{p}(m)-S_{p}(p^{2}m)}{p-1}=\frac{(p^{2}-1)S_{p}(m)}{p-1}=(p+1)S_{p}(m)\geq p+1.$
Hence (3.5) gives
$\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)-2\geq p-1>0,$
contradicting the assumption that the valuation is 0. Thus $d(e)\geq 1$.
Step 3: from $\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=0$ we obtain $c(e)=d(e)=1$. By Step 1, whenever some occurring positive index is not divisible by $p$ we have $c(e)\geq 1$, and by Step 2 we always have $d(e)\geq 1$. Since (3.5) says
$0=\operatorname{ord}_{p}\bigl(\gamma_{B}(e)\bigr)=c(e)+d(e)-2,$
we must have
$c(e)=d(e)=1.$
Step 4: $d(e)=1$ forces every $e_{n}$ to be divisible by $p$. The condition $d(e)=1$ means that
$\sum_{n=0}^{k-1}S_{p}(e_{n})-S_{p}\!\left(\sum_{n=0}^{k-1}e_{n}\right)=p-1,$
so the addition of the integers $e_{0},e_{1},\dots,e_{k-1}$ to obtain $q=p^{2}$ produces exactly one carry in base $p$. Now suppose some $e_{n}$ has nonzero units digit. Then in the sum
$\sum_{n=0}^{k-1}e_{n}=p^{2},$
there is a carry from the units place to the $p$-place. But the final result $p^{2}$ has base-$p$ digits $(0,0,1)$, so in order for the $p$-place digit also to end up equal to 0, there must be a second carry from the $p$-place to the $p^{2}$-place. This contradicts the fact that there is only one carry. Hence every $e_{n}$ is divisible by $p$.
So we may write
$e_{n}=pf_{n},\qquad\sum_{n=0}^{k-1}f_{n}=p.$
Then
$k=\sum_{n=1}^{k-1}ne_{n}=p\sum_{n=1}^{k-1}f_{n}n,$
so
$S_{p}(k)=S_{p}\!\left(\sum_{n=1}^{k-1}f_{n}n\right)\leq\sum_{n=1}^{k-1}f_{n}S_{p}(n).$
Therefore
$c(e)=\frac{p\sum_{n=1}^{k-1}f_{n}S_{p}(n)-S_{p}(k)}{p-1}\geq\sum_{n=1}^{k-1}f_{n}S_{p}(n).$
Since $c(e)=1$, it follows that
$\sum_{n=1}^{k-1}f_{n}S_{p}(n)\leq 1.$
The left-hand side is a sum of nonnegative integers, and it is positive because $k>0$; thus exactly one positive index $n$ satisfies $f_{n}=1$, and for that index we have $S_{p}(n)=1$. Therefore $n$ is a power of $p$, and the positive part of the monomial is $a_{n}^{p}$.
If this $n$ is not divisible by $p$, then the only possibility is $n=1$, and since $\sum f_{n}=p$ we must have $f_{0}=p-1$, that is,
$e_{0}=p(p-1)=q-p,\qquad e_{1}=p.$
This gives the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$. In all other cases, every positive index that occurs is divisible by $p$. ∎
Corollary 3.4 (Dependence only on the earlier $a_{pi}$).
If $p\mid k$ and $k>p$, then modulo $p$ the quantity $a_{k}$ depends only on the earlier coefficients $a_{pi}$ with $pi<k$.
Proof.
By Propositions 2.1 and 2.2,
$a_{k}\equiv-B_{k}[x^{k}]\pmod{p}\qquad(p\mid k).$
Now Proposition 3.3 shows that any monomial in $B_{k}[x^{k}]$ whose coefficient can survive modulo $p$ is either the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$ or else involves only indices divisible by $p$. If $k>p$, then the exceptional monomial cannot occur, because its weight is
$0\cdot(q-p)+1\cdot p=p,$
not $k$. Therefore every surviving monomial in $B_{k}[x^{k}]$ involves only $a_{pi}$ with $pi<k$. ∎
Lemma 3.5.
The first nontrivial coefficient of $f_{\phi}$ is
$a_{1}=-1.$
Proof.
Write
$f_{\phi}(x)=x+a_{1}x^{2}+O(x^{3}).$
Then
$\phi(f_{\phi}(x))=f_{\phi}(x)^{q}+p^{2}f_{\phi}(x)^{q+1}.$
Expanding to order $x^{q+1}$ gives
$f_{\phi}(x)^{q}=x^{q}+qa_{1}x^{q+1}+O(x^{q+2}),\qquad p^{2}f_{\phi}(x)^{q+1}=p^{2}x^{q+1}+O(x^{q+2}).$
On the other hand,
$f_{\phi}(x^{q})=x^{q}+O(x^{2q}),$
so comparison of the $x^{q+1}$-coefficients in the Böttcher equation yields
$qa_{1}+p^{2}=0.$
Since $q=p^{2}$, we obtain $a_{1}=-1$. ∎
Lemma 3.6.
One has
$a_{p}\equiv-1\pmod{p}$.
Proof.
By Propositions 2.1 and 2.2,
$a_{p}\equiv-B_{p}[x^{p}]\pmod{p}$.
Since there are no positive indices less than $p$ that are divisible by $p$, Proposition 3.3 shows that the only monomial in $B_{p}[x^{p}]$ whose coefficient can be a $p$-adic unit is the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$. Its coefficient is
$\gamma_{p}=\frac{p!}{q}\binom{q}{q-p,p}=\frac{1}{p^{2}}\prod_{j=0}^{p-1}(p^{2}-j).$
Reducing modulo $p$ gives
$\gamma_{p}\equiv\prod_{j=1}^{p-1}(-j)=(-1)^{p-1}(p-1)!\equiv-1\pmod{p}$
by Wilson’s theorem. Using $a_{0}=1$ and Lemma 3.5, we find
$B_{p}[x^{p}]\equiv\gamma_{p}a_{1}^{p}\equiv(-1)(-1)^{p}\equiv 1\pmod{p}$.
Therefore
$a_{p}\equiv-1\pmod{p}$.
Proof of Theorem 2.3.
We argue by induction on $M$. The case $M=1$ is Lemma 3.6.
Assume now that $M\geq 2$ and that
(3.6) $a_{pi}\equiv(-1)^{i}\pmod{p}\qquad(1\leq i<M).$
Applying Corollary 3.4 with $k=pM$, we obtain
(3.7) $a_{pM}\equiv-B_{pM}[x^{pM}]\pmod{p}$,
and only the indices divisible by $p$ can contribute. Therefore, after setting $y=x^{p}$,
(3.8) $B_{pM}[x^{pM}]\equiv\frac{(pM)!}{q}[y^{M}]\left(\sum_{i=0}^{M-1}\frac{a_{pi}}{(pi)!}y^{i}\right)^{q}\pmod{p}$.
Here $a_{0}=1$ corresponds to $i=0$.
For $i\geq 0$, define
$\nu_{i}:=\frac{(pi)!}{p^{i}i!}.$
Then $\nu_{i}\in\mathbf{Z}$ and
(3.9) $\nu_{i}=\prod_{r=1}^{i}\prod_{t=1}^{p-1}\bigl(p(r-1)+t\bigr)\equiv((p-1)!)^{i}\equiv(-1)^{i}\pmod{p}$
by Wilson’s theorem. Hence, by the induction hypothesis (3.6),
$\omega_{i}:=a_{pi}\nu_{i}^{-1}\in\mathbf{Z}_{(p)}$
satisfies $\omega_{i}\equiv 1\pmod{p}$ for $0\leq i<M$ (with $\omega_{0}=1$). Consequently,
(3.10) $\frac{a_{pi}}{(pi)!}=\frac{\omega_{i}}{p^{i}i!}\qquad(0\leq i<M).$
Write $\omega_{i}=1+pz_{i}$ with $z_{i}\in\mathbf{Z}_{(p)}$, and define
$H_{M}(y):=\sum_{i=0}^{M-1}\frac{y^{i}}{p^{i}i!},\qquad L_{M}(y):=\sum_{i=0}^{M-1}z_{i}\frac{y^{i}}{p^{i}i!}.$
Then by (3.10),
$\sum_{i=0}^{M-1}\frac{a_{pi}}{(pi)!}y^{i}=H_{M}(y)+pL_{M}(y).$
We claim that
(3.11) $\frac{p^{M}M!}{q}[y^{M}]\bigl(H_{M}(y)+pL_{M}(y)\bigr)^{q}\equiv\frac{p^{M}M!}{q}[y^{M}]H_{M}(y)^{q}\pmod{p}$.
Indeed,
$\bigl(H_{M}+pL_{M}\bigr)^{q}-H_{M}^{q}=\sum_{j=1}^{q}\binom{q}{j}p^{j}H_{M}^{q-j}L_{M}^{j}.$
Any term contributing to the coefficient of $y^{M}$ in $H_{M}^{q-j}L_{M}^{j}$ has total denominator $p^{M}\prod i!^{e_{i}}$, so
$y^{M}]H_{M}^{q-j}L_{M}^{j}\in p^{-M}\mathbf{Z}_{(p)}.$
Therefore it is enough to show that
$\operatorname{ord}_{p}\!\left(\frac{p^{j}}{q}\binom{q}{j}\right)\geq 1\qquad(1\leq j\leq q).$
Using
$\binom{q}{j}=\frac{q}{j}\binom{q-1}{j-1}$,
we get
$\frac{p^{j}}{q}\binom{q}{j}=\frac{p^{j}}{j}\binom{q-1}{j-1}$,
so
$\operatorname{ord}_{p}\!\left(\frac{p^{j}}{q}\binom{q}{j}\right)\geq j-\operatorname{ord}_{p}(j)\geq 1.$
This proves (3.11).
Now combine (3.7), (3.8), and (3.11). Since $(pM)!=p^{M}M!\nu_{M}$, we obtain
(3.12) $a_{pM}\equiv-\nu_{M}\frac{p^{M}M!}{q}[y^{M}]H_{M}(y)^{q}\pmod{p}$.
By Corollary 3.2,
$y^{M}]H_{M}(y)^{q}=\frac{p^{M}}{M!}-\frac{q}{p^{M}M!}.$
Substituting into (3.12) gives
$a_{pM}\equiv-\nu_{M}\frac{p^{M}M!}{q}\left(\frac{p^{M}}{M!}-\frac{q}{p^{M}M!}\right)\equiv-\nu_{M}\bigl(p^{2M-2}-1\bigr)\pmod{p}$.
Since $M\geq 2$, we have $p^{2M-2}\equiv 0\pmod{p}$, so
$a_{pM}\equiv\nu_{M}\pmod{p}$.
Finally, (3.9) gives $\nu_{M}\equiv(-1)^{M}\pmod{p}$, and hence
$a_{pM}\equiv(-1)^{M}\pmod{p}$.
This completes the induction. ∎
4. Vector partitions and the $B$-term for $a\neq 0$
The higher base-$p$ digits of the index will be encoded by vectors.
Definition 4.1.
Fix $r\geq 1$. A digit vector is an element
$d=(d_{1},\dots,d_{r})\in\{0,1,\dots,p-1\}^{r}.$
Its weight is
$|d|:=d_{1}+\cdots+d_{r},$
and its numeric value is
$N(d):=\sum_{i=1}^{r}d_{i}p^{i}.$
A vector partition of $d$ is a formal product
$\Pi=\prod_{\beta\neq 0}\beta^{r_{\beta}},$
with $r_{\beta}\in\mathbf{Z}_{\geq 0}$ and
$\sum_{\beta\neq 0}r_{\beta}\beta=d.$
The total number of blocks is
$R(\Pi):=\sum_{\beta\neq 0}r_{\beta}.$
Fix $a\in\{0,1,\dots,p-1\}$. For a digit vector $\beta$ define
$H_{\beta}(y):=\sum_{j=0}^{a}\frac{a_{N(\beta)+j}}{j!}y^{j},\qquad F_{a}(y):=\sum_{j=0}^{a}\frac{a_{j}}{j!}y^{j},\qquad\widetilde{H}_{d}(y):=\sum_{j=0}^{a-1}\frac{a_{N(d)+j}}{j!}y^{j}.$
Write also
$d!:=\prod_{i=1}^{r}d_{i}!,\qquad\beta!:=\prod_{i=1}^{r}\beta_{i}!\qquad\text{for }\beta=(\beta_{1},\dots,\beta_{r}).$
Proposition 4.2 (Vector-partition expansion of the $B$-term).
Let $k=N(d)+a$ with $1\leq a\leq p-1$. Then
(4.1) $B_{k}[x^{k}]\equiv a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}+\sum_{\begin{subarray}{c}\Pi\vdash d\\ \Pi\neq(d)\end{subarray}}(-1)^{R(\Pi)-1}(R(\Pi)-1)!\frac{d!}{\prod_{\beta\neq 0}r_{\beta}!(\beta!)^{r_{\beta}}}\prod_{\beta\neq 0}\left(\frac{H_{\beta}(y)}{F_{a}(y)}\right)^{r_{\beta}}\right)\pmod{p}$.
Here $(d)$ denotes the one-block partition of $d$.
Proof.
Let $a^{e}=\prod_{n\geq 0}a_{n}^{e_{n}}$ be a monomial occurring in $B_{k}[x^{k}]$. Then
$\sum_{n\geq 0}e_{n}=p^{2},\qquad\sum_{n\geq 0}ne_{n}=k.$
By (2.5),
$\operatorname{ord}_{p}\bigl(B_{k}[x^{k}][a^{e}]\bigr)=c(e)+d(e)-2,$
where $c(e),d(e)\in\mathbf{Z}_{\geq 0}$ measure the carry defects in the additions
$\sum_{n}e_{n}n=k,\qquad\sum_{n}e_{n}=p^{2}.$
Hence a monomial can survive modulo $p$ only if
$c(e)+d(e)=2.$
We first show that, because $k\equiv a\not\equiv 0\pmod{p}$, one must in fact have
$d(e)=2,\qquad c(e)=0.$
Indeed, if $d(e)=0$, then $\sum_{n}S_{p}(e_{n})=1$. Writing $e_{n}=pu_{n}+r_{n}$ with $0\leq r_{n}\leq p-1$, we have
$\sum_{n}r_{n}\equiv\sum_{n}e_{n}=p^{2}\equiv 0\pmod{p}$.
If some $r_{n}$ were nonzero, then $\sum_{n}r_{n}\geq p$, contradicting $\sum_{n}S_{p}(e_{n})=1$. Thus every $e_{n}$ is divisible by $p$. Dividing by $p$ and repeating the same argument shows that every $e_{n}$ is divisible by $p^{2}$, so exactly one $e_{t}$ equals $p^{2}$. Then $k=p^{2}t$, contradicting $p\nmid k$.
If $d(e)=1$, then $\sum_{n}S_{p}(e_{n})=p$. Again write $e_{n}=pu_{n}+r_{n}$ with $0\leq r_{n}\leq p-1$. Since $\sum_{n}e_{n}=p^{2}$, the sum $\sum_{n}r_{n}$ is a multiple of $p$. If some $r_{n}$ were nonzero, then $\sum_{n}r_{n}\geq p$; because
$\sum_{n}S_{p}(e_{n})=\sum_{n}r_{n}+\sum_{n}S_{p}(u_{n})=p,$
this would force $\sum_{n}r_{n}=p$ and $\sum_{n}S_{p}(u_{n})=0$, hence every $u_{n}=0$ and $\sum_{n}e_{n}=p$, impossible. Therefore all $r_{n}$ vanish, so every $e_{n}$ is divisible by $p$. But then $k=\sum_{n}e_{n}n$ is divisible by $p$, again a contradiction. This proves that $d(e)=2$, hence $c(e)=0$.
Because $c(e)=0$, the addition $\sum_{n}e_{n}n=k$ is carry-free in base $p$. Write each positive index uniquely as
$n=N(\beta)+j,\qquad 0\leq j\leq a,$
where $\beta=(\beta_{1},\dots,\beta_{r})$ records the digits in positions $p,p^{2},\dots,p^{r}$. For $\beta\neq 0$ set
$m_{\beta,j}:=e_{N(\beta)+j},\qquad r_{\beta}:=\sum_{j=0}^{a}m_{\beta,j},$
and for the zero vector set
$n_{j}:=e_{j}\qquad(1\leq j\leq a).$
The carry-free condition is equivalent to
$\sum_{\beta\neq 0}r_{\beta}\beta=d,\qquad\sum_{\beta\neq 0}\sum_{j=0}^{a}jm_{\beta,j}+\sum_{j=1}^{a}jn_{j}=a.$
Thus the nonzero vectors occurring form a vector partition
$\Pi=\prod_{\beta\neq 0}\beta^{r_{\beta}}\qquad\text{of }d.$
The only excluded index is $k=N(d)+a$, so when $\Pi=(d)$ the choice $m_{d,a}=1$ and all other positive multiplicities zero must be removed.
Fix such a partition $\Pi$, and set
$R:=R(\Pi)=\sum_{\beta\neq 0}r_{\beta},\qquad t:=\sum_{j=1}^{a}n_{j},\qquad M:=R+t.$
Then $e_{0}=p^{2}-M$, and the coefficient of the corresponding monomial is
$B_{k}[x^{k}][a^{e}]$ $=k!\,\frac{(p^{2}-1)!}{(p^{2}-M)!\,\prod_{\beta,j}m_{\beta,j}!\,\prod_{j=1}^{a}n_{j}!}\prod_{\beta,j}\frac{1}{(N(\beta)+j)!^{m_{\beta,j}}}\prod_{j=1}^{a}\frac{1}{j!^{n_{j}}}.$
Modulo $p$,
$\frac{(p^{2}-1)!}{(p^{2}-M)!}=\prod_{u=1}^{M-1}(p^{2}-u)\equiv(-1)^{M-1}(M-1)!.$
Also, because the addition of the positive indices is carry-free, Lucas’ theorem in multinomial form gives
$\frac{k!}{\prod_{\beta,j}(N(\beta)+j)!^{m_{\beta,j}}\prod_{j=1}^{a}j!^{n_{j}}}\equiv\frac{a!\,d!}{\prod_{\beta,j}j!^{m_{\beta,j}}(\beta!)^{m_{\beta,j}}\prod_{j=1}^{a}j!^{n_{j}}}\pmod{p}$.
Hence
$B_{k}[x^{k}][a^{e}]$ $\equiv(-1)^{M-1}(M-1)!\,\frac{a!\,d!}{\prod_{\beta,j}m_{\beta,j}!\,j!^{m_{\beta,j}}(\beta!)^{m_{\beta,j}}\,\prod_{j=1}^{a}n_{j}!\,j!^{n_{j}}}\prod_{\beta,j}a_{N(\beta)+j}^{m_{\beta,j}}\prod_{j=1}^{a}a_{j}^{n_{j}}.$
Now sum first over the zero-vector blocks. Since
$F_{a}(y)-1=\sum_{j=1}^{a}\frac{a_{j}}{j!}y^{j},$
we have the negative-binomial identity
$\frac{1}{F_{a}(y)^{R}}=\sum_{t\geq 0}(-1)^{t}\binom{R+t-1}{t}\bigl(F_{a}(y)-1\bigr)^{t}.$
After multiplying by (R1)!(R-1)!, this exactly reproduces the factor
$(-1)^{t}(R+t-1)!$
coming from the zero-vector blocks. Next, for each fixed nonzero $\beta$,
$\frac{1}{r_{\beta}!}\bigl(H_{\beta}(y)\bigr)^{r_{\beta}}=\sum_{\sum_{j}m_{\beta,j}=r_{\beta}}\prod_{j=0}^{a}\frac{1}{m_{\beta,j}!}\left(\frac{a_{N(\beta)+j}}{j!}y^{j}\right)^{m_{\beta,j}}.$
Collecting all $\beta$ therefore yields the contribution of all monomials with fixed vector partition $\Pi$:
$(-1)^{R-1}(R-1)!\,\frac{d!}{\prod_{\beta\neq 0}r_{\beta}!(\beta!)^{r_{\beta}}}\prod_{\beta\neq 0}\left(\frac{H_{\beta}(y)}{F_{a}(y)}\right)^{r_{\beta}}.$
Finally, when $\Pi=(d)$ one must delete the forbidden monomial corresponding to the excluded index $k=N(d)+a$; this is exactly accomplished by replacing $H_{d}(y)$ with the truncated polynomial
$\widetilde{H}_{d}(y)=\sum_{j=0}^{a-1}\frac{a_{N(d)+j}}{j!}y^{j}.$
Extracting the coefficient of $y^{a}$ now gives (4.1). ∎
5. Low coefficients and a multivariate cumulant collapse for $a\neq 0$
The first block of coefficients admits a particularly simple description modulo $p$.
Proposition 5.1 (The first block modulo $p$).
For $0\leq n\leq p-1$ one has
$a_{n}\equiv(-1)^{n}(n+1)^{n-1}\pmod{p}$,
with the convention that the case $n=0$ gives $a_{0}=1$.
Proof.
Let
$g(x):=\sum_{n\geq 0}\frac{a_{n}}{n!}x^{n},$
so that $f_{\phi}(x)=xg(x)$ and $g(0)=1$. The Böttcher equation becomes
$g(x^{q})=g(x)^{q}\bigl(1+p^{2}xg(x)\bigr),\qquad q=p^{2}.$
Taking formal logarithms gives
(5.1) $\log g(x^{q})=q\log g(x)+\log\bigl(1+p^{2}xg(x)\bigr).$
Fix $1\leq n\leq p-1$. Since $q=p^{2}>n$, the coefficient of $x^{n}$ on the left-hand side of (5.1) is 0. Expanding the second logarithm, we obtain
$\log\bigl(1+p^{2}xg(x)\bigr)=p^{2}xg(x)+\sum_{m\geq 2}(-1)^{m-1}\frac{p^{2m}}{m}x^{m}g(x)^{m}.$
Hence the $x^{n}$-coefficient of the second term is congruent modulo $p^{3}$ to the $x^{n}$-coefficient of $p^{2}xg(x)$, because every summand with $m\geq 2$ carries the factor $p^{2m}$ with $2m\geq 4$. Dividing the coefficient relation by $p^{2}$ therefore yields
$x^{n}]\log g(x)\equiv-[x^{n-1}]g(x)\pmod{p}$.
Since $g(x)-F_{p-1}(x)=O(x^{p})$, we conclude that
(5.2) $\log F_{p-1}(x)\equiv-xF_{p-1}(x)\pmod{(p,x^{p})}$.
Set
$U(x):=xF_{p-1}(x)\in x+x^{2}\mathbf{F}_{p}[x]/(x^{p+1}).$
Exponentiating (5.2) in $\mathbf{F}_{p}[x]/(x^{p})$ gives
$F_{p-1}(x)\equiv e^{-xF_{p-1}(x)}\pmod{(p,x^{p})}$,
and therefore
$U(x)=xF_{p-1}(x)\equiv xe^{-U(x)}\pmod{(p,x^{p+1})}$.
Equivalently,
$U(x)e^{U(x)}\equiv x\pmod{(p,x^{p+1})}$.
By Lagrange inversion, the unique formal solution of $U=xe^{-U}$ satisfies
$x^{m}]U(x)=\frac{1}{m}[u^{m-1}]e^{-mu}=\frac{(-m)^{m-1}}{m!}\qquad(m\geq 1).$
Thus, modulo $(p,x^{p+1})$,
$U(x)\equiv\sum_{m=1}^{p}\frac{(-m)^{m-1}}{m!}x^{m}.$
Dividing by $x$ gives
$F_{p-1}(x)\equiv 1+\sum_{n=1}^{p-1}\frac{(-(n+1))^{n}}{(n+1)!}x^{n}\equiv 1+\sum_{n=1}^{p-1}(-1)^{n}(n+1)^{n-1}\frac{x^{n}}{n!}\pmod{(p,x^{p})}$.
Comparing coefficients proves the proposition. ∎
Corollary 5.2.
For every $1\leq a\leq p-1$ one has
$\log F_{a}(y)\equiv-yF_{a}(y)\pmod{(p,y^{a+1})}$,
and hence
$a![y^{a}]\log F_{a}(y)\equiv-a\,a_{a-1}\pmod{p}$.
Proof.
Since $F_{a}(y)-F_{p-1}(y)=O(y^{a+1})$, Proposition 5.1 and (5.2) give
$\log F_{a}(y)\equiv\log F_{p-1}(y)\equiv-yF_{p-1}(y)\equiv-yF_{a}(y)\pmod{(p,y^{a+1})}$.
Taking the coefficient of $y^{a}$ yields
$y^{a}]\log F_{a}(y)\equiv-[y^{a-1}]F_{a}(y)=-\frac{a_{a-1}}{(a-1)!}\pmod{p}$,
and multiplying by $a!$ gives the stated formula. ∎
The expansion (4.1) becomes tractable because, after the inductive substitutions, every block polynomial depends only on the total weight of the corresponding digit vector.
Lemma 5.3 (Induction-to-block transfer).
Assume that (1.1) is already known for every index of the form
$N(\beta)+j\qquad(0\leq j\leq a,\ \beta\neq 0,\ |\beta|<s).$
Then for every nonzero digit vector $\beta$ with $|\beta|<s$ one has
$H_{\beta}(y)\equiv(-D)^{|\beta|}F_{a}(y)\pmod{p},\qquad D:=1+y\frac{d}{dy}.$
Consequently, for every digit vector $d$ of weight $s$,
$\widetilde{H}_{d}(y)\equiv(-D)^{s}F_{a}(y)+(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}\pmod{p}$.
Proof.
Fix a nonzero digit vector $\beta$ with $|\beta|<s$. By the induction hypothesis from (1.1), for every $0\leq j\leq a$ we have
$a_{N(\beta)+j}\equiv(-1)^{|\beta|}(j+1)^{|\beta|}a_{j}\pmod{p}$.
Summing over $j$ gives
$H_{\beta}(y)=\sum_{j=0}^{a}\frac{a_{N(\beta)+j}}{j!}y^{j}\equiv\sum_{j=0}^{a}(-1)^{|\beta|}(j+1)^{|\beta|}\frac{a_{j}}{j!}y^{j}=(-D)^{|\beta|}F_{a}(y)\pmod{p}$,
because
$D^{m}F_{a}(y)=\sum_{j=0}^{a}(j+1)^{m}\frac{a_{j}}{j!}y^{j}\qquad(m\geq 0).$
This proves the first assertion.
Now let $d$ be a digit vector of weight $s$. For $0\leq j\leq a-1$, the same induction hypothesis gives
$a_{N(d)+j}\equiv(-1)^{s}(j+1)^{s}a_{j}\pmod{p}$.
Therefore
$\widetilde{H}_{d}(y)\equiv\sum_{j=0}^{a-1}(-1)^{s}(j+1)^{s}\frac{a_{j}}{j!}y^{j}\pmod{p}$.
Comparing with
$D^{s}F_{a}(y)=\sum_{j=0}^{a}(j+1)^{s}\frac{a_{j}}{j!}y^{j},$
we see that the only missing term is the top coefficient $j=a$, namely
$(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}.$
Hence
$\widetilde{H}_{d}(y)\equiv(-D)^{s}F_{a}(y)+(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!}\pmod{p}$.
This proves the lemma. ∎
For nonzero digit vectors define
$U_{\beta}(y):=(-1)^{|\beta|}\frac{D^{|\beta|}F_{a}(y)}{F_{a}(y)}.$
Introduce formal variables $t=(t_{1},\dots,t_{r})$ and set
$t^{\beta}:=\prod_{i=1}^{r}t_{i}^{\beta_{i}}\qquad\text{for }\beta=(\beta_{1},\dots,\beta_{r}).$
The associated multivariate moment series is
$M(t):=1+\sum_{\beta\neq 0}U_{\beta}(y)\frac{t^{\beta}}{\beta!}.$
Define the multivariate cumulants $K_{d}(U)$ by
(5.3) $\log M(t)=\sum_{d\neq 0}K_{d}(U)\frac{t^{d}}{d!}.$
Proposition 5.4 (Collapse to one variable).
Let $d=(d_{1},\dots,d_{r})$ be a nonzero digit vector and set $s:=|d|$. Then
$a![y^{a}]K_{d}(U)=(-1)^{s}a![y^{a}]N^{s}\log F_{a}(y),\qquad N:=y\frac{d}{dy}.$
Consequently,
(5.4) $a![y^{a}]K_{d}(U)\equiv(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
Proof.
Because $U_{\beta}(y)$ depends only on $|\beta|$, the multivariate series depends only on
$T:=t_{1}+\cdots+t_{r}.$
Indeed,
$\sum_{|\beta|=m}\frac{t^{\beta}}{\beta!}=\frac{T^{m}}{m!},$
so
$M(t)$ $=1+\sum_{m\geq 1}(-1)^{m}\frac{D^{m}F_{a}(y)}{F_{a}(y)}\frac{T^{m}}{m!}$
$=\frac{\exp(-TD)F_{a}(y)}{F_{a}(y)}$
$=\frac{e^{-T}F_{a}(ye^{-T})}{F_{a}(y)}.$
Here we used $D=1+N$ and the identity $e^{-TN}G(y)=G(ye^{-T})$ for every polynomial $G$. Taking logarithms gives
$\log M(t)=-T+\log F_{a}(ye^{-T})-\log F_{a}(y).$
Expanding the middle term as a Taylor series in the operator $N$ yields
$\log F_{a}(ye^{-T})=\sum_{m\geq 0}(-1)^{m}N^{m}\log F_{a}(y)\frac{T^{m}}{m!}.$
Since the coefficient of $t^{d}/d!$ in a series depending only on $T$ is the coefficient of $T^{|d|}/|d|!$, we obtain
$a![y^{a}]K_{d}(U)=(-1)^{s}a![y^{a}]N^{s}\log F_{a}(y).$
The linear term T-T contributes only when $s=1$, and even then its $y^{a}$-coefficient is 0 because $a\geq 1$.
Finally,
$a![y^{a}]N^{s}\log F_{a}(y)=a^{s}\,a![y^{a}]\log F_{a}(y),$
so Corollary 5.2 gives
$a![y^{a}]K_{d}(U)=(-1)^{s}a^{s}\,a![y^{a}]\log F_{a}(y)\equiv(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
This proves the proposition. ∎
6. Proof of (1.1)
We now prove (1.1). The case $a=0$ is already known from Theorem 2.3, so it remains to treat the residue classes $1\leq a\leq p-1$.
Proof of (1.1).
We argue by outer induction on $a\in\{0,1,\dots,p-1\}$. The case $a=0$ is Theorem 2.3. Fix $1\leq a\leq p-1$ and assume that the theorem is already known for every smaller residue class $0,1,\dots,a-1$.
For this fixed $a$, we argue by inner induction on
$s=d_{1}+\cdots+d_{r}.$
If $s=0$, then $k=a$ and the desired congruence is tautological.
Now let $s\geq 1$, let $d=(d_{1},\dots,d_{r})$ be a digit vector of weight $s$, and write
$k=N(d)+a.$
By Proposition 2.2 and the already established formula in residue class a1a-1 (for $a=1$ this is Theorem 2.3), we have
(6.1) $C_{k}[x^{k}]\equiv a\,a_{k-1}\equiv(-1)^{s}a^{s+1}a_{a-1}\pmod{p}$.
By Proposition 2.1,
(6.2) $A_{k}[x^{k}]\equiv 0\pmod{p}$.
For the $B$-term, set
$V_{\beta}(y):=\frac{H_{\beta}(y)}{F_{a}(y)}\qquad(\beta\neq 0).$
Let $K_{d}(V)$ be defined by the same cumulant formula as in (5.3), with $U_{\beta}$ replaced by $V_{\beta}$. Since the one-block partition contributes exactly $V_{d}(y)=H_{d}(y)/F_{a}(y)$, Proposition 4.2 may be rewritten as
(6.3) $B_{k}[x^{k}]\equiv a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}+K_{d}(V)-V_{d}(y)\right)\pmod{p}$.
Now every block appearing in a proper partition of $d$ has weight strictly smaller than $s$. Therefore Lemma 5.3, applied with the inner induction hypothesis from (1.1), gives
$V_{\beta}(y)=\frac{H_{\beta}(y)}{F_{a}(y)}\equiv(-1)^{|\beta|}\frac{D^{|\beta|}F_{a}(y)}{F_{a}(y)}=U_{\beta}(y)\pmod{p}$
for every such $\beta$. Therefore the proper-partition contribution is unchanged if we replace $V_{\beta}$ by $U_{\beta}$:
(6.4) $a![y^{a}]\bigl(K_{d}(V)-V_{d}(y)\bigr)\equiv a![y^{a}]\bigl(K_{d}(U)-U_{d}(y)\bigr)\pmod{p}$.
On the other hand, Lemma 5.3 gives
$\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}-U_{d}(y)\equiv(-1)^{s-1}(a+1)^{s}a_{a}\frac{y^{a}}{a!F_{a}(y)}\pmod{p}$.
Since $F_{a}(y)=1+O(y)$, we get
(6.5) $a![y^{a}]\left(\frac{\widetilde{H}_{d}(y)}{F_{a}(y)}-U_{d}(y)\right)=(-1)^{s-1}(a+1)^{s}a_{a}.$
Combining (6.3), (6.4), (6.5), and Proposition 5.4, we obtain
(6.6) $B_{k}[x^{k}]\equiv(-1)^{s-1}(a+1)^{s}a_{a}+(-1)^{s+1}a^{s+1}a_{a-1}\pmod{p}$.
Finally, substitute (6.2), (6.6), and (6.1) into the recursion (2.1):
$a_{k}$ $\equiv-\left((-1)^{s-1}(a+1)^{s}a_{a}+(-1)^{s+1}a^{s+1}a_{a-1}\right)-(-1)^{s}a^{s+1}a_{a-1}\pmod{p}$
$=(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
This completes both inductions and proves the theorem. ∎
7. Proof of (1.2) and Theorem 1.2
We first promote Proposition 5.1 to the fully explicit form of Theorem $1.$
Proof of (1.2).
Let $k\geq 1$ and write
$k=a+d_{1}p+d_{2}p^{2}+\cdots+d_{r}p^{r},\qquad 0\leq a,d_{i}\leq p-1,$
with $s=S_{p}(k)-a$.
If $a=0$, then part (i) gives
$a_{k}\equiv(-1)^{s}\pmod{p}$.
Since $S_{p}(k)=s$ and $a+1=1$, this is exactly the desired formula.
Assume now that $1\leq a\leq p-1$. By part (i),
$a_{k}\equiv(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
Applying Proposition 5.1 with $n=a$ gives
$a_{a}\equiv(-1)^{a}(a+1)^{a-1}\pmod{p}$.
Substituting this into the previous congruence yields
$a_{k}\equiv(-1)^{s+a}(a+1)^{s+a-1}\pmod{p}$.
Since $s+a=S_{p}(k)$, this is exactly the asserted formula. ∎
We now deduce the original residue-class conjecture of Salerno and Silverman.
Proof of Theorem 1.2.
The first congruence
$a_{pm}\equiv(-1)^{m}\pmod{p}$
is exactly Theorem 2.3.
For the second congruence, write
pm1=(m1)p+(p1).pm-1=(m-1)p+(p-1).
If $m=1$, Proposition 5.1 gives $a_{p-1}\equiv 0\pmod{p}$. If $m\geq 2$, then (1.1) yields
$a_{pm-1}\equiv(-1)^{m-1}p^{m-1}a_{p-1}\equiv 0\pmod{p}$.
Thus $a_{pm-1}\equiv 0\pmod{p}$ for every $m\geq 1$.
For the third congruence, write
pm2=(m1)p+(p2).pm-2=(m-1)p+(p-2).
Then (1.1) gives
$a_{pm-2}\equiv(-1)^{m-1}(p-1)^{m-1}a_{p-2}\equiv a_{p-2}\pmod{p}$.
By Proposition 5.1,
$a_{p-2}\equiv(-1)^{p-2}(p-1)^{p-3}\equiv-1\pmod{p}$.
Therefore
$a_{pm-2}\equiv-1\pmod{p}$
for every $m\geq 1$. ∎

### 06 Apr 2026 v2

@@
-A Digit-Sum Formula for Böttcher Coordinates
+Digit-sum Formulas and Divisible Classes for Böttcher Coordinates of $x^{p^{2}}+p^{r+2}x^{p^{2}+1}$
 Abstract.
-Let $p$ be an odd prime and let $\phi(x)=x^{p^{2}}+p^{2}x^{p^{2}+1}$.
-We prove an unconditional digit-sum congruence formula for the coefficients of the Böttcher coordinate of $\phi$.
-The argument separates the divisible case $k\equiv 0\pmod{p}$ from the general case, gives a complete vector-partition description of the surviving $B$-term monomials, and reduces the remaining partition sum to a one-variable cumulant.
-As consequences we recover the residue-class congruences conjectured by Salerno and Silverman and obtain an explicit description of the first block of coefficients modulo $p$.
+Let $p$ be an odd prime and, for $r\geq 0$, let
+$\phi_{r}(x)=x^{p^{2}}+p^{r+2}x^{p^{2}+1}.$
+For $r=0$ we prove a complete digit-sum congruence formula modulo $p$ for the coefficients of the Böttcher coordinate of $\phi_{0}$.
+As consequences we prove Conjecture 25 of Salerno and Silverman and obtain an explicit description of the first block of coefficients modulo $p$.
+For $r\geq 1$ we prove on the divisible non-pure class a leading term theorem together with a global digit-weight lower bound.
+These results yield an unconditional proof of Conjecture 27(a) of Salerno and Silverman.
 Key words and phrases:
@@
 1. Introduction
-The local study of holomorphic iteration begins with the functional equations of Schröder and Koenigs, and the theory of complex dynamics took its modern shape through the early twentieth-century work of Böttcher, Fatou, and Julia [1, 6].
-For a one-dimensional analytic germ with a superattracting fixed point, Böttcher’s theorem provides a canonical local coordinate in which the map becomes a pure monomial.
-In polynomial dynamics this coordinate at infinity does considerably more than linearize the map: it identifies the basin of infinity with the exterior disk, encodes the escape-rate function, and underlies external rays, landing theory, and the combinatorial study of Julia sets and the Mandelbrot set [2, 3, 6].
-The same normal form also appears in several variables, where the local geometry near superattracting fixed points becomes subtler [5].
-In recent years Böttcher coordinates have also become important in non-archimedean dynamics.
-Salerno and Silverman studied their integrality properties for one-dimensional superattracting germs over $p$-adic rings [8], while Fu and Nie later settled a conjecture of Salerno–Silverman on the radius of convergence in the wild superattracting case [4].
+The local study of superattracting germs goes back to Böttcher’s theorem, which provides a canonical local coordinate in which a germ becomes a pure monomial.
+In complex dynamics this coordinate plays a central role both near a superattracting fixed point and, for polynomials, at infinity, where it encodes the escape-rate function and the external combinatorics of Julia sets and the Mandelbrot set; see, for example, [1, 2, 3, 6, 7].
+The same normal form also appears in higher-dimensional holomorphic dynamics [5].
+In the non-archimedean setting, Böttcher coordinates have an additional arithmetic aspect.
+Salerno and Silverman studied their integrality properties for one-dimensional superattracting germs and formulated several precise conjectures in the wild case [8].
+More recently, Fu and Nie proved the expected radius of convergence in a general wild superattracting setting [4]; in particular, their theorem settles Conjecture 27(b) for the family considered here.
 The family relevant for the present paper is
-$\phi(x)=x^{p^{2}}+p^{2}x^{p^{2}+1}.$
-If
-$f_{\phi}(x)=x\sum_{k\geq 0}\frac{a_{k}}{k!}x^{k}$
+$\phi_{r}(x)=x^{p^{2}}+p^{r+2}x^{p^{2}+1}\qquad(r\geq 0),$
+where $p\geq 3$ is an odd prime. If
+$f_{r}(x)=x\sum_{k\geq 0}\frac{a_{k}(r)}{k!}x^{k}$
 denotes the Böttcher coordinate characterized by
-$\phi\bigl(f_{\phi}(x)\bigr)=f_{\phi}(x^{p^{2}}),$
-then Salerno and Silverman recorded striking mod-$p$ patterns in the coefficients $a_{k}$; their original formulation is Conjecture 25 of [8].
-For an integer $n=\sum_{i\geq 0}n_{i}p^{i}$ with $0\leq n_{i}\leq p-1$, let
-$S_{p}(n):=\sum_{i\geq 0}n_{i}$
-be the sum of its base-$p$ digits.
-From now on, we assume that $p\geq 3$.
-We note that the corresponding result for $p=2$ can be established similarly, but since the arguments for all $p\geq 3$ fit into a unified framework, we do not treat the case $p=2$ in the present paper.
-The following are our main theorems.
-Theorem 1.1 (Digit-sum formula).
+$\phi_{r}\bigl(f_{r}(x)\bigr)=f_{r}(x^{p^{2}}),$
+then Conjectures 25 and 27 of [8] predict striking congruence and valuation patterns for the coefficients $a_{k}(r)$.
+Since Conjecture 27(b) is already supplied by Fu and Nie, the higher-fiber part of the present paper is concerned only with Conjecture 27(a).
+The special fiber $r=0$ already exhibits a rich mod-$p$ structure.
+In that case one expects a digit-sum formula for every coefficient, not only on the divisible class $p\mid k$ but on every residue class modulo $p$.
+Our first result establishes the full congruence law.
+Theorem 1.1 (Digit-sum formula in the special fiber).
 Let $k\geq 1$ and write
@@
 with
-$s:=d_{1}+\cdots+d_{r}=S_{p}(k)-a.$
+$s:=d_{1}+\cdots+d_{r}.$
 Then
 (1.1) $a_{k}\equiv(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
-In fact, we have a stronger result
-(1.2) $a_{k}\equiv(-1)^{S_{p}(k)}(a+1)^{S_{p}(k)-1}\pmod{p}$.
-We also recover the original residue-class conjecture of Salerno and Silverman.
-Theorem 1.2 (Salerno–Silverman’s Conjecture 25).
+In fact, one has the stronger closed formula
+(1.2) $a_{k}\equiv(-1)^{a+s}(a+1)^{a+s-1}\pmod{p}$.
+As an immediate consequence we recover the residue-class congruences conjectured in [8, Conjecture 25].
+Theorem 1.2 (Salerno–Silverman’s Conjecture 25 for $r=0$).
 For every integer $m\geq 1$ one has
 $a_{pm}\equiv(-1)^{m},\qquad a_{pm-1}\equiv 0,\qquad a_{pm-2}\equiv-1\pmod{p}$.
-The proof is organized by separating the residue class $a=0$ from the classes $a\neq 0$.
-Section 2 recalls the AABBCC recursion and proves the basic $A$- and $C$-term lemmas.
-Section 3 supplies a complete proof of the divisible case $a=0$.
-The remaining sections treat $a\neq 0$: Section 4 gives a vector-partition expansion of the $B$-term, Section 5 evaluates the first block of coefficients and collapses the multivariate cumulant, and Section 6 proves (1.1).
-In the final section we prove (1.2) and deduce Theorem 1.2.
+For the higher fibers $r\geq 1$, the relevant structure is valuation-theoretic. If
+$v_{i}(r):=\operatorname{ord}_{p}\bigl(a_{p^{i}}(r)\bigr)\qquad(i\geq 0),$
+and if
+$m=m_{0}+m_{1}p+\cdots+m_{s}p^{s},\qquad 0\leq m_{i}\leq p-1,$
+we set
+$\Lambda_{m}(r):=\sum_{i=0}^{s}m_{i}v_{i+1}(r),\qquad\epsilon_{m}:=m_{0}+\cdots+m_{s}=S_{p}(m).$
+Our third main theorem gives the precise leading term on the divisible non-pure class and, from it, proves the valuation statement Conjecture 27(a) of [8].
+Theorem 1.3 (Divisible classes and Conjecture 27(a) for $r\geq 1$).
+Fix $r\geq 1$. Then the following hold.
+  1. (a)
+    If $m\geq 1$ is not a power of $p$, then
+    $a_{pm}(r)=(-1)^{\epsilon_{m}}p^{\Lambda_{m}(r)}+O\bigl(p^{\Lambda_{m}(r)+1}\bigr),$
+    where
+    $v_{0}(r)=r,\qquad v_{1}(r)=pr,$
+    and, for every $n\geq 2$,
+    $v_{n}(r)=\begin{cases}(p+1)p^{n-2}-2+v_{n-2}(r),&\text{if $n$ is even and $r\geq n-1$,}\\[3.0pt] p\,v_{n-1}(r),&\text{otherwise.}\end{cases}$
+    In particular,
+    $\operatorname{ord}_{p}\bigl(a_{pm}(r)\bigr)=\Lambda_{m}(r).$
+  2. (b)
+    If $p\mid k$, then
+    $\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)=\frac{1-p^{-r}}{p-1}k+O(1).$
+The special-fiber part of the paper is driven by a complete analysis of the AABBCC recursion modulo $p$, together with a vector-partition description of the surviving $B$-term monomials.
+The higher-fiber part is organized around a simultaneous induction on three main assertions—a global lower bound, a leading theorem on the divisible non-pure class, and a pure-power leading theorem—together with the resulting subadditivity of the digit weight.
+Once these ingredients are in place, the pure-power branch pattern and Conjecture 27(a) follow.
+We do not treat the case $p=2$ here. The arguments for odd $p$ already exhibit the full mechanism, and keeping $p\geq 3$ avoids a number of low-characteristic notational distractions.
 2. The recursion and two structural lemmas
+In this part we work in the special fiber $r=0$, so $\phi=\phi_{0}$ and $a_{k}=a_{k}(0)$.
 Let $q:=p^{2}$. As in [8, §5], write
@@
 $A_{k}:=\frac{k!}{q}\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{q\ell},\qquad B_{k}:=\frac{k!}{q}\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{q},\qquad C_{k}:=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}}{\ell!}x^{\ell}\right)^{q+1}.$
-Here $P[x^{n}]$ denotes the coefficient of $x^{n}$ in the power series $P(x)$. Throughout we use Legendre’s formula
+Here $P[x^{n}]$ denotes the coefficient of $x^{n}$ in the power series $P(x)$. For a nonnegative integer
+$N=\sum_{i\geq 0}N_{i}p^{i},\qquad 0\leq N_{i}\leq p-1,$
+we write
+$S_{p}(N):=\sum_{i\geq 0}N_{i},\qquad T_{p}(N):=\frac{N!}{p^{\operatorname{ord}_{p}(N!)}},$
+so $T_{p}(N)$ is the prime-to-$p$ part of $N!$. Throughout we use Legendre’s formula
 (2.2) $\operatorname{ord}_{p}(n!)=\frac{n-S_{p}(n)}{p-1}$
@@
 $U(x)e^{U(x)}\equiv x\pmod{(p,x^{p+1})}$.
-By Lagrange inversion, the unique formal solution of $U=xe^{-U}$ satisfies
+By Lagrange inversion (see [9, 10]), the unique formal solution of $U=xe^{-U}$ satisfies
 $x^{m}]U(x)=\frac{1}{m}[u^{m-1}]e^{-mu}=\frac{(-m)^{m-1}}{m!}\qquad(m\geq 1).$
@@
 This proves the proposition. ∎
-6. Proof of (1.1)
+6. Proof of Theorems 1.1 and 1.2
+Proof of (1.1)
 We now prove (1.1). The case $a=0$ is already known from Theorem 2.3, so it remains to treat the residue classes $1\leq a\leq p-1$.
@@
 This completes both inductions and proves the theorem. ∎
-7. Proof of (1.2) and Theorem 1.2
-We first promote Proposition 5.1 to the fully explicit form of Theorem $1.$
+Proof of (1.2) and Theorem 1.2
+We first combine Proposition 5.1 with (1.1) to obtain the closed formula in Theorem $1.$
 Proof of (1.2).
@@
 with $s=S_{p}(k)-a$.
-If $a=0$, then part (i) gives
+If $a=0$, then (1.1) gives
 $a_{k}\equiv(-1)^{s}\pmod{p}$.
 Since $S_{p}(k)=s$ and $a+1=1$, this is exactly the desired formula.
-Assume now that $1\leq a\leq p-1$. By part (i),
+Assume now that $1\leq a\leq p-1$. By (1.1),
 $a_{k}\equiv(-1)^{s}(a+1)^{s}a_{a}\pmod{p}$.
@@
 for every $m\geq 1$. ∎
+7. The family $\phi_{r}$ for $r\geq 1$: recursion and coefficient valuations
+Fix a prime $p$ and set $q:=p^{2}$. Let
+$\phi_{r}(x)=x^{q}+p^{r+2}x^{q+1}=x^{q}\bigl(1+qp^{r}x\bigr).$
+The Böttcher equation
+$\phi_{r}\bigl(f_{r}(x)\bigr)=f_{r}(x^{q})$
+gives the usual recursion from [8, §4, §7]. Write
+$f_{r}(x)=x\sum_{k\geq 0}\frac{a_{k}(r)}{k!}x^{k},\qquad a_{0}(r)=1,$
+and define
+(7.1) $A_{k}(r):=\frac{k!}{q}\sum_{\ell=0}^{k-1}\frac{a_{\ell}(r)}{\ell!}x^{q\ell},\qquad B_{k}(r):=\frac{k!}{q}\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}(r)}{\ell!}x^{\ell}\right)^{q},\qquad C_{k}(r):=k!x\left(\sum_{\ell=0}^{k-1}\frac{a_{\ell}(r)}{\ell!}x^{\ell}\right)^{q+1}.$
+Then
+(7.2) $a_{k}(r)=A_{k}(r)[x^{k}]-B_{k}(r)[x^{k}]-p^{r}C_{k}(r)[x^{k}].$
+This is exactly the same formal decomposition as (2.1); only the outer factor $p^{r}$ is new.
+Since $r$ is fixed throughout the remainder of the paper, we write $A_{k},B_{k},C_{k}$ in place of $A_{k}(r),B_{k}(r),C_{k}(r)$ whenever no confusion can arise.
+Recall the Legendre formula
+(7.3) $\operatorname{ord}_{p}(N!)=\frac{N-S_{p}(N)}{p-1}$
+for the sum-of-digits function $S_{p}(N)$.
+The $A$-term
+If $q\nmid k$ then $A_{k}[x^{k}]=0$. If $k=qm$, then
+(7.4) $A_{k}[x^{k}]=\frac{(qm)!}{q\,m!}a_{m}.$
+Since $S_{p}(qm)=S_{p}(m)$, Legendre’s formula gives
+(7.5) $\operatorname{ord}_{p}\!\left(\frac{(qm)!}{q\,m!}\right)=(p+1)m-2.$
+The $B$- and $C$-terms
+The combinatorial coefficient formulas for $B_{k}[x^{k}]$ and $C_{k}[x^{k}]$ are identical to those in Sections 2 and 3, since the parameter $r$ enters only through the outer factor $p^{r}$ in (7.2).
+Accordingly, the valuation formulas (3.4) and (3.5) remain valid verbatim for the coefficients of $B_{k}[x^{k}]$.
+In particular, all $B$-coefficients are $p$-integral, and when $p\mid k$ every unit-coefficient monomial is governed by Proposition 3.3: either it is the exceptional monomial $a_{0}^{q-p}a_{1}^{p}$, or else all of its positive indices are divisible by $p$.
+For the $C$-term we use the coefficient formula proved in Proposition 2.2.
+In the present family this gives the same divisibility statement as before: if $p\mid k$, then every coefficient of $C_{k}[x^{k}]$ is divisible by $p$.
+We record this as
+(7.6) $\operatorname{ord}_{p}\bigl(C_{k}[x^{k}][a^{e}]\bigr)\geq 1\qquad(p\mid k)$
+for every monomial $a^{e}$ occurring in $C_{k}[x^{k}]$.
+8. The simultaneous induction for $r\geq 1$
+For $n\geq 0$ put
+$v_{n}(r):=\operatorname{ord}_{p}\bigl(a_{p^{n}}(r)\bigr).$
+When $r$ is fixed we often suppress it from the notation and write simply $v_{n}$. Thus $v_{0}(r)=r$ because comparing the coefficient of $x^{q+1}$ in the Böttcher relation gives
+(8.1) $a_{1}(r)=-p^{r}.$
+Definition 8.1.
+If $k=\sum_{i\geq 0}k_{i}p^{i}$ is the base-$p$ expansion of $k$, define its digit weight by
+(8.2) $\Lambda_{r}(k):=\sum_{i\geq 0}k_{i}v_{i}(r)$
+and its digit monomial by
+(8.3) $\mathfrak{m}_{r}(k):=\prod_{i\geq 0}a_{p^{i}}(r)^{k_{i}}.$
+Then
+$\operatorname{ord}_{p}\bigl(\mathfrak{m}_{r}(k)\bigr)=\Lambda_{r}(k).$
+Proposition 8.2 (Exact two-term decomposition on pure powers).
+For every $n\geq 2$ there exist polynomials
+$R_{n},S_{n}\in\mathbf{Z}_{p}[a_{1}(r),\dots,a_{p^{n}-1}(r)]$
+such that
+(8.4) $a_{p^{n}}(r)=\alpha_{n}a_{p^{n-2}}(r)-\gamma_{n}a_{p^{n-1}}(r)^{p}+pR_{n}+p^{r+1}S_{n},$
+where
+(8.5) $\alpha_{n}:=\frac{(p^{n})!}{q\,(p^{n-2})!},\qquad\operatorname{ord}_{p}(\alpha_{n})=(p+1)p^{n-2}-2,\qquad p^{-\operatorname{ord}_{p}(\alpha_{n})}\alpha_{n}\equiv 1\pmod{p}$,
+and
+(8.6) $\gamma_{n}\in\mathbf{Z}_{p}^{\times},\qquad\gamma_{n}\equiv-1\pmod{p}$.
+Proof.
+When $k=p^{n}=qp^{n-2}$, only the term $\ell=p^{n-2}$ contributes to $A_{k}[x^{k}]$, so
+$A_{k}[x^{k}]=\frac{(p^{n})!}{q\,(p^{n-2})!}a_{p^{n-2}}=\alpha_{n}a_{p^{n-2}}.$
+The valuation statement for $\alpha_{n}$ is exactly (7.5).
+Its unit congruence follows from Legendre’s formula together with the prime-to-$p$ factorial congruence $T_{p}(p^{m})\equiv(-1)^{m}\pmod{p}$ used in [8, Lemma 22(a)].
+On the $B$-side, Theorem 5 of [8] shows that in $B_{p^{n}}[x^{p^{n}}]$ the unique monomial whose coefficient is a $p$-adic unit is the exceptional monomial $a_{0}^{q-p}a_{p^{n-1}}^{p}$.
+Its coefficient is
+$\gamma_{n}=\frac{(p^{n})!}{q}\binom{q}{q-p,p}\frac{1}{(p^{n-1}!)^{p}}=\frac{(p^{n})!}{p\,(p^{n-1}!)^{p}}\binom{p^{2}-1}{p-1}$.
+Since
+$\operatorname{ord}_{p}\!\left(\frac{(p^{n})!}{(p^{n-1}!)^{p}}\right)=1,$
+the first factor is a $p$-adic unit. Writing it in terms of prime-to-$p$ parts gives
+$\frac{(p^{n})!}{p\,(p^{n-1}!)^{p}}=\frac{T_{p}(p^{n})}{T_{p}(p^{n-1})^{p}}\equiv\frac{(-1)^{n}}{((-1)^{n-1})^{p}}\equiv-1\pmod{p}$,
+because $p$ is odd. Moreover Lucas’s theorem gives
+$\binom{p^{2}-1}{p-1}\equiv\binom{p-1}{0}\binom{p-1}{p-1}\equiv 1\pmod{p}$.
+Hence $\gamma_{n}\equiv-1\pmod{p}$.
+All remaining $B$-monomials have coefficients divisible by $p$, and this is exactly the polynomial $pR_{n}$.
+Finally, because $p\mid p^{n}$, every coefficient of $C_{p^{n}}[x^{p^{n}}]$ is divisible by $p$ by (7.6).
+Multiplying by the outer factor $p^{r}$ in (7.2) gives the term $p^{r+1}S_{n}$.
+∎
+We now organize the argument for $r\geq 1$ as a simultaneous strong induction. We consider the following recursive assertions:
+Statement 8.3 (Global digit-weight lower bound).
+For every positive integer $k$, one has
+$\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)\geq\Lambda_{r}(k).$
+Statement 8.4 (Leading term on the divisible non-pure class).
+If $p\mid k$ and $k$ is not a power of $p$, then
+(8.7) $a_{k}(r)\equiv\mathfrak{m}_{r}(k)\pmod{p^{\Lambda_{r}(k)+1}}$.
+Statement 8.5 (Pure-power leading theorem).
+For every integer $n\geq 2$, define
+(8.8) $A_{n}(r):=(p+1)p^{n-2}-2+v_{n-2}(r),\qquad B_{n}(r):=p\,v_{n-1}(r).$
+Then
+(8.9) $v_{n}(r)=\min\{A_{n}(r),B_{n}(r)\},$
+and, whenever $A_{n}(r)\neq B_{n}(r)$,
+(8.10) $a_{p^{n}}(r)=\begin{cases}\alpha_{n}a_{p^{n-2}}(r)+O\bigl(p^{v_{n}(r)+1}\bigr),&\text{if }A_{n}(r)<B_{n}(r),\\[3.0pt] -\gamma_{n}a_{p^{n-1}}(r)^{p}+O\bigl(p^{v_{n}(r)+1}\bigr),&\text{if }B_{n}(r)<A_{n}(r).\end{cases}$
+In particular,
+$v_{n}(r)\leq p\,v_{n-1}(r)\qquad(n\geq 1).$
+Statement 8.6 (Subadditivity of the digit weight).
+For all $m,n\geq 0$ one has
+$\Lambda_{r}(m+n)\leq\Lambda_{r}(m)+\Lambda_{r}(n).$
+More generally, for every finite sum $N=n_{1}+\cdots+n_{t}$,
+$\Lambda_{r}(N)\leq\Lambda_{r}(n_{1})+\cdots+\Lambda_{r}(n_{t}).$
+To keep the induction parameters explicit, we proceed as follows.
+For each $n\geq 2$ we first prove Statement 8.5 at the pure-power index $p^{n}$ from Statements 8.3 and 8.4 below $p^{n}$.
+Once Statement 8.5 is known up to level $n$, the inequalities $v_{j}(r)\leq p\,v_{j-1}(r)$ for $1\leq j\leq n$ imply Statement 8.6 for every integer whose base-$p$ expansion uses only the digits $1,p,\dots,p^{n}$.
+We then prove Statement 8.4 at each divisible non-pure index $k$, and finally Statement 8.3 at the general index $k$.
+Proof of Statement 8.5 at level $n$ and Statement 8.6 up to level $n$.
+Fix $n\geq 2$.
+Assume Statement 8.3 and Statement 8.4 hold for every $k<p^{n}$, and Statement 8.5 holds for every level $2\leq j<n$.
+By the induction scheme explained above, Statement 8.6 is therefore already available for every integer whose highest base-$p$ digit is at most $p^{n-1}$.
+In particular,
+$v_{j}(r)\leq p\,v_{j-1}(r)\qquad(1\leq j\leq n-1),$
+because this is part of Statement 8.5 for $j\geq 2$, while for $j=1$ it is simply $v_{1}(r)=pr=pv_{0}(r)$.
+We first control the term $pR_{n}$.
+The polynomial $R_{n}$ is the sum of the non-exceptional monomials occurring in $B_{p^{n}}[x^{p^{n}}]$.
+If the coefficient of such a monomial is divisible by $p$, then the outer factor $p$ in $pR_{n}$ already contributes one extra $p$.
+If the coefficient is a $p$-adic unit, then Proposition 3.3 forces every positive index in that monomial to be divisible by $p$.
+For each lower divisible non-pure index we may then replace $a_{m}(r)$ by its leading digit monomial using Statement 8.4; on the pure-power line the same replacement is tautological from the definition of $\mathfrak{m}_{r}$.
+After this substitution, the only carry-free projected monomial of weight $B_{n}(r)=p\,v_{n-1}(r)$ is the exceptional monomial $a_{p^{n-1}}(r)^{p}$, and that monomial has been removed from $R_{n}$.
+Hence every term in $pR_{n}$ has valuation strictly larger than $B_{n}(r)$, and therefore strictly larger than $\min\{A_{n}(r),B_{n}(r)\}$.
+For $p^{r+1}S_{n}$, every coefficient of $C_{p^{n}}[x^{p^{n}}]$ is divisible by $p$ by (7.6).
+Every monomial in $S_{n}$ is a product of lower coefficients whose total index is $p^{n}-1$.
+Hence Statement 8.3 for lower indices and Statement 8.6 at lower levels show that every monomial contribution to $p^{r+1}S_{n}$ has valuation at least
+$r+1+\Lambda_{r}(p^{n}-1).$
+Now $\Lambda_{r}(p^{n}-1)=(p-1)\sum_{i=0}^{n-1}v_{i}(r)$, because the base-$p$ expansion of $p^{n}-1$ has all digits equal to p1p-1 below level $n$.
+Using $v_{j+1}(r)\leq p\,v_{j}(r)$ for $0\leq j\leq n-2$, we obtain
+$\sum_{i=0}^{n-2}(p-1)v_{i}(r)\geq\sum_{i=0}^{n-2}\bigl(v_{i+1}(r)-v_{i}(r)\bigr)=v_{n-1}(r)-v_{0}(r).$
+Therefore
+$r+1+\Lambda_{r}(p^{n}-1)$ $=1+v_{0}(r)+(p-1)\sum_{i=0}^{n-1}v_{i}(r)$
+$\geq 1+v_{0}(r)+v_{n-1}(r)-v_{0}(r)+(p-1)v_{n-1}(r)$
+$=1+p\,v_{n-1}(r)=1+B_{n}(r).$
+Hence every contribution in $p^{r+1}S_{n}$ has valuation strictly larger than $B_{n}(r)$, and therefore strictly larger than $\min\{A_{n}(r),B_{n}(r)\}$.
+Substituting these bounds into Proposition 8.2 gives (8.9) and, whenever $A_{n}(r)\neq B_{n}(r)$, the leading branch formula (8.10).
+The displayed inequality $v_{n}(r)\leq p\,v_{n-1}(r)$ follows immediately from (8.9).
+To deduce Statement 8.6 up to level $n$, let $N=n_{1}+\cdots+n_{t}$ be a finite sum whose highest base-$p$ digit occurs at level at most $n$.
+Passing from the separate base-$p$ expansions of the $n_{i}$ to the base-$p$ expansion of $N$ amounts to replacing carries of $p$ copies of $p^{j}$ by one copy of $p^{j+1}$.
+Since $v_{j+1}(r)\leq p\,v_{j}(r)$ for every $j\leq n-1$, each carry can only decrease (or preserve) the total weight.
+Iterating over all carries yields
+$\Lambda_{r}(N)\leq\Lambda_{r}(n_{1})+\cdots+\Lambda_{r}(n_{t}),$
+which proves Statement 8.6 for all integers supported in the first $n+1$ digit places. ∎
+Proof of Statement 8.4 at the index $k=pm$.
+Let $k=pm$ with $m\geq 1$ not a power of $p$.
+Assume Statement 8.3 and Statement 8.4 hold for every $n^{\prime}<k$, and Statement 8.5 holds for every pure-power level $p^{j}<k$.
+Then Statement 8.6 is available for all integers $\leq k$.
+Claim 8.7 (Shift inequality).
+For every integer $m^{\prime}\geq 0$ with $p^{2}m^{\prime}\leq k$ one has
+$\Lambda_{r}(p^{2}m^{\prime})\leq(p+1)m^{\prime}-2+\Lambda_{r}(m^{\prime}).$
+If $m^{\prime}$ is not a power of $p$, then the inequality is strict.
+Proof of the claim.
+Write $m^{\prime}=\sum_{i}m^{\prime}_{i}p^{i}$. Since Statement 8.5 has already been proved at every level appearing in the base-$p$ expansion of $m^{\prime}$, we have
+$v_{i+2}(r)\leq(p+1)p^{i}-2+v_{i}(r)$
+for every $i$ with $m^{\prime}_{i}>0$. Therefore
+$\Lambda_{r}(p^{2}m^{\prime})=\sum_{i}m^{\prime}_{i}v_{i+2}(r)\leq\sum_{i}m^{\prime}_{i}\bigl((p+1)p^{i}-2+v_{i}(r)\bigr)=(p+1)m^{\prime}-2\sum_{i}m^{\prime}_{i}+\Lambda_{r}(m^{\prime}).$
+Since $\sum_{i}m^{\prime}_{i}\geq 1$, this proves the first assertion. If $m^{\prime}$ is not a power of $p$, then $\sum_{i}m^{\prime}_{i}\geq 2$, so the inequality is strict. ∎
+If $q\nmid k$, then $A_{k}[x^{k}]=0$. If $k=qm^{\prime}$, then $m^{\prime}$ is not a power of $p$, so Claim 8.7 is strict. Hence
+$\operatorname{ord}_{p}\bigl(A_{k}[x^{k}]\bigr)\geq(p+1)m^{\prime}-2+\Lambda_{r}(m^{\prime})>\Lambda_{r}(k).$
+For the $C$-term, every monomial coefficient is divisible by $p$ because $p\mid k$; see (7.6). Therefore
+$\operatorname{ord}_{p}\bigl(p^{r}C_{k}[x^{k}]\bigr)\geq 1+r+\Lambda_{r}(k-1)>\Lambda_{r}(k)$
+by Statement 8.6. Thus
+(8.11) $a_{k}(r)\equiv-B_{k}[x^{k}]\pmod{p^{\Lambda_{r}(k)+1}}$.
+If a monomial in $B_{k}[x^{k}]$ contains some positive index not divisible by $p$, then either its coefficient is divisible by $p$ or, by Proposition 3.3, it cannot be a unit-coefficient monomial on the class $p\mid k$.
+In either case Statement 8.3 shows that its total valuation is $\Lambda_{r}(k)$.
+Hence modulo $p^{\Lambda_{r}(k)+1}$ one may keep only the divisible truncation and write
+(8.12) $B_{k}[x^{k}]\equiv\frac{k!}{q}[y^{m}]\left(\sum_{i=0}^{m-1}\frac{a_{pi}(r)}{(pi)!}y^{i}\right)^{q}\pmod{p^{\Lambda_{r}(k)+1}}$.
+For every $i<m$, either $pi$ is a pure power, in which case $a_{pi}(r)=\mathfrak{m}_{r}(pi)$ by definition, or $pi$ is divisible and non-pure, in which case the induction hypothesis gives
+$a_{pi}(r)\equiv\mathfrak{m}_{r}(pi)\pmod{p^{\Lambda_{r}(pi)+1}}$.
+Because the coefficient in front of each monomial in (8.12) is $p$-integral, every occurrence of one of the error terms raises the total valuation by at least one. Therefore
+(8.13) $B_{k}[x^{k}]\equiv\frac{k!}{q}[y^{m}]\left(\sum_{i=0}^{m-1}\frac{\mathfrak{m}_{r}(pi)}{(pi)!}y^{i}\right)^{q}\pmod{p^{\Lambda_{r}(k)+1}}$.
+Set
+$H_{m}(y):=\sum_{i=0}^{m-1}\frac{\mathfrak{m}_{r}(pi)}{(pi)!}y^{i}.$
+We now determine the coefficient of the monomial $\mathfrak{m}_{r}(k)$ in $\frac{k!}{q}[y^{m}]H_{m}(y)^{q}$ modulo $p$. After projecting to $\mathfrak{m}_{r}(k)$, only carry-free tails survive. Write
+$m=m_{0}+m_{1}p+\cdots+m_{s}p^{s},\qquad 0\leq m_{\nu}\leq p-1.$
+Running the coefficient extraction digit by digit, exactly as in the handwritten digit argument, one reduces the scalar coefficient to successive one-digit problems.
+At each nonterminal digit level one encounters the full truncated exponential coming from the allowable tails, and Corollary 3.2 shows that its normalized contribution is congruent to $1\pmod{p}$.
+At the terminal level—the highest nonzero digit of mm—one gets the proper truncation, and the same one-digit calculation as in the base case gives contribution $1\pmod{p}$.
+Therefore the total coefficient of $\mathfrak{m}_{r}(k)$ in $\frac{k!}{q}[y^{m}]H_{m}(y)^{q}$ is congruent to $1\pmod{p}$.
+Equivalently,
+$B_{k}[x^{k}]\equiv-\mathfrak{m}_{r}(k)\pmod{p^{\Lambda_{r}(k)+1}}$.
+Combining this with (8.11) and (8.13) yields
+$a_{k}(r)\equiv\mathfrak{m}_{r}(k)\pmod{p^{\Lambda_{r}(k)+1}}$,
+which is Statement 8.4 at the index $k$. ∎
+Proof of Statement 8.3 at the index $k$.
+Fix $k\geq 1$.
+Assume Statement 8.3 holds for every $n^{\prime}<k$, Statement 8.4 holds for every divisible non-pure $n^{\prime}<k$, and Statement 8.5 holds for every pure-power level $p^{j}\leq k$.
+Then Statement 8.6 is available for all integers $\leq k$.
+If $k=p^{n}$ is a pure power, then Statement 8.5 gives
+$\operatorname{ord}_{p}\bigl(a_{p^{n}}(r)\bigr)=v_{n}(r)=\Lambda_{r}(p^{n}),$
+so the lower bound holds with equality. We may therefore assume that $k$ is not a pure power.
+We use the recursion (7.2).
+The AA-term. If $q\nmid k$, then $A_{k}[x^{k}]=0$. If $k=qm$, then by (7.4), the induction hypothesis, and Claim 8.7,
+$\operatorname{ord}_{p}\bigl(A_{k}[x^{k}]\bigr)\geq(p+1)m-2+\Lambda_{r}(m)\geq\Lambda_{r}(qm)=\Lambda_{r}(k).$
+The $B$-term. Take a monomial $\gamma_{B}(e)a^{e}$ in $B_{k}[x^{k}]$. By (3.5), $\gamma_{B}(e)\in\mathbf{Z}_{p}$. By the induction hypothesis,
+$\operatorname{ord}_{p}(a^{e})\geq\sum_{n=0}^{k-1}e_{n}\Lambda_{r}(n).$
+Since $\sum_{n=0}^{k-1}e_{n}n=k$, Statement 8.6 gives
+$\sum_{n=0}^{k-1}e_{n}\Lambda_{r}(n)\geq\Lambda_{r}(k).$
+Hence every $B$-monomial has valuation at least $\Lambda_{r}(k)$, so
+$\operatorname{ord}_{p}\bigl(B_{k}[x^{k}]\bigr)\geq\Lambda_{r}(k).$
+The $C$-term. Take a monomial $\gamma_{C}(e)a^{e}$ in $C_{k}[x^{k}]$. By the coefficient formula in Proposition 2.2, the coefficient $\gamma_{C}(e)$ is $p$-integral, and by the induction hypothesis,
+$\operatorname{ord}_{p}(a^{e})\geq\sum_{n=0}^{k-1}e_{n}\Lambda_{r}(n)\geq\Lambda_{r}(k-1)$
+because $\sum_{n=0}^{k-1}e_{n}n=k-1$. Therefore
+$\operatorname{ord}_{p}\bigl(p^{r}\gamma_{C}(e)a^{e}\bigr)\geq r+\Lambda_{r}(k-1)\geq\Lambda_{r}(k),$
+where the last inequality is the special case k=(k1)+1k=(k-1)+1 of Statement 8.6. Since every summand in (7.2) has valuation at least $\Lambda_{r}(k)$, so does $a_{k}(r)$. ∎
+The pure-power branch pattern and Conjecture 27(a)
+We now return to the pure powers and deduce the branch pattern and the asymptotic law predicted in Conjecture 27(a) of [8]. For $n\geq 2$, Statement 8.5 reads
+$v_{n}(r)=\min\{A_{n}(r),B_{n}(r)\},\qquad A_{n}(r):=(p+1)p^{n-2}-2+v_{n-2}(r),\qquad B_{n}(r):=p\,v_{n-1}(r).$
+Again we suppress the argument $(r)$ whenever no confusion can arise, and we set $s:=\left\lfloor\frac{r+1}{2}\right\rfloor$.
+Proposition 8.8 (Pure-power branch pattern).
+For every $j\geq 1$ one has:
+  1. (1)
+    every odd layer $n=2j+1$ is $B$-dominated;
+  2. (2)
+    an even layer $n=2j$ is $A$-dominated if and only if $r\geq 2j-1$;
+  3. (3)
+    once $n\geq 2s+1$, all subsequent layers are $B$-dominated.
+Proof.
+The proof is now unconditional because Statement 8.5 has already been proved.
+The initial values are
+$v_{0}=r,\qquad v_{1}=pr,\qquad v_{2}=r+p-1,\qquad v_{3}=p(r+p-1).$
+Assume that $2j$ is $A$-dominated, so $v_{2j}=A_{2j}$. Then the previous odd layer is $B$-dominated, hence $v_{2j-1}=pv_{2j-2}$. Therefore
+$A_{2j+1}-B_{2j+1}$ $=\bigl((p+1)p^{2j-1}-2\bigr)+v_{2j-1}-pv_{2j}$
+$=\bigl((p+1)p^{2j-1}-2\bigr)+pv_{2j-2}-p\bigl((p+1)p^{2j-2}-2+v_{2j-2}\bigr)$
+$=2p-2>0,$
+so every odd layer is $B$-dominated.
+Now assume that up to level 2j22j-2 the pattern is alternating $B,A,B,A,\dots,B,A$. Then
+$v_{2j-2}=r+\sum_{m=1}^{j-1}\bigl((p+1)p^{2m-2}-2\bigr)=r-2(j-1)+\frac{p^{2j-2}-1}{p-1}.$
+Hence
+$A_{2j}-B_{2j}$ $=\bigl((p+1)p^{2j-2}-2\bigr)+v_{2j-2}-p^{2}v_{2j-2}$
+$=-(p-1)\bigl((p+1)(r-2j)+2p+1\bigr).$
+Thus $A_{2j}<B_{2j}$ if and only if $r\geq 2j-1$. This proves the second assertion.
+Finally, if $n\geq 2s+1$ is odd, then it is $B$-dominated by part (1).
+If $n\geq 2s+2$ is even, write $n=2j$.
+Then $j\geq s+1$, so r<2j1r<2j-1 by the definition of $s$, and part (2) shows that $n$ is $B$-dominated.
+This proves part (3).
+Proposition 8.9 (Normalized pure-power units).
+For every $n\geq 0$ one has
+$p^{-v_{n}(r)}a_{p^{n}}(r)\equiv-1\pmod{p}$.
+Proof.
+We argue by induction on $n$.
+The case $n=0$ is (8.1).
+For $n=1$, one has $A_{p}[x^{p}]=0$, while $p^{r}C_{p}[x^{p}]$ has valuation at least $pr+1$ by (7.6).
+The unique $B$-monomial of valuation $pr$ is the exceptional monomial $a_{0}^{q-p}a_{1}(r)^{p}$, and its coefficient is the analogous unit $\gamma_{1}$ obtained by the same calculation as in the proof of Proposition 8.2; in particular $\gamma_{1}\equiv-1\pmod{p}$.
+Hence
+$a_{p}(r)\equiv-\gamma_{1}a_{1}(r)^{p}\equiv-p^{pr}\pmod{p^{pr+1}}$,
+so the claim holds for $n=1$.
+Now let $n\geq 2$ and assume the claim for all smaller indices.
+By Proposition 8.8, the layer $n$ is either $A$-dominated or $B$-dominated, so Statement 8.5 applies with one of the branches in (8.10).
+If $A_{n}(r)<B_{n}(r)$, then
+$p^{-v_{n}(r)}a_{p^{n}}(r)\equiv p^{-A_{n}(r)}\alpha_{n}a_{p^{n-2}}(r)\equiv-1\pmod{p}$,
+because $p^{-\operatorname{ord}_{p}(\alpha_{n})}\alpha_{n}\equiv 1\pmod{p}$ and the induction hypothesis gives the unit of $a_{p^{n-2}}(r)$.
+If $B_{n}(r)<A_{n}(r)$, then (8.6) and the induction hypothesis yield
+$p^{-v_{n}(r)}a_{p^{n}}(r)\equiv p^{-B_{n}(r)}\bigl(-\gamma_{n}a_{p^{n-1}}(r)^{p}\bigr)\equiv(-\gamma_{n})\bigl(p^{-v_{n-1}(r)}a_{p^{n-1}}(r)\bigr)^{p}\equiv 1\cdot(-1)^{p}\equiv-1\pmod{p}$.
+Thus the claim holds for every $n\geq 0$. ∎
+Corollary 8.10 (Stable pure-power slope).
+For every $n\geq 2s$ one has
+$v_{n}(r)=\frac{1-p^{-r}}{p-1}p^{n}.$
+Proof.
+At level $2s$ one computes directly from the closed form in the proof of Proposition 8.8 that
+$v_{2s}(r)=\frac{1-p^{-r}}{p-1}p^{2s}.$
+All higher levels are $B$-dominated, so $v_{n}(r)=p^{n-2s}v_{2s}(r)$. ∎
+We can now prove Conjecture 27(a).
+Theorem 8.11 (Conjecture 27(a)).
+Let $k\geq 1$ satisfy $p\mid k$. Then
+$\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)=\frac{1-p^{-r}}{p-1}k+O(1).$
+More precisely, if $k=\sum_{i\geq 0}k_{i}p^{i}$, then
+$\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)=\sum_{i\geq 0}k_{i}v_{i}(r),$
+and the error with respect to the linear function $\frac{1-p^{-r}}{p-1}k$ is uniformly bounded.
+Proof.
+If $k$ is a power of $p$, this is Corollary 8.10. Otherwise $p\mid k$ and Statement 8.4 gives
+$\operatorname{ord}_{p}(a_{k}(r))=\Lambda_{r}(k)=\sum_{i\geq 0}k_{i}v_{i}(r).$
+For all sufficiently large $i$, namely $i\geq 2s$, Corollary 8.10 gives
+$v_{i}(r)=\frac{1-p^{-r}}{p-1}p^{i}.$
+Therefore the difference
+$\Lambda_{r}(k)-\frac{1-p^{-r}}{p-1}k$
+comes only from the finitely many digits below level $2s$, hence is bounded independently of $k$. ∎
+The radius statement Conjecture 27(b) is already available from the general theorem of Fu and Nie [4], so the higher-fiber analysis in the present paper is complete at this point.